A nearsighted person cannot see beyond 25 cm. Assuming that the separation of the glass from the eye is 1 cm, find the power of lens needed to see distant objects.

Grade:upto college level

1 Answers

Deepak Patra
9 years ago
Sol. For the near sighted person, v = distance of image from glass = distance of image from eye – separation between glass and eye = 25 cm – 1cm = 24 cm = 0.24m So, for the glass, u = ∞ and v = – 24 cm = –0.24m So, 1/f = 1/v – 1/u = 1/-0.24 – 1/∞ = - 4.2 D